The length of rectangular garden is 5 meters more than its width.If the perimeter is 130 meters,find its dimensions

Answer by JulietG(1812)

The perimeter is a fence around. It's 2 widths + 2 lengths.
130 = 2(L) + 2(W)
L = W+5
Substitute the value of L from the bottom equation into the top.
130 = 2(W+5) + 2W
Distribute
130 = 2W + 10 + 2W
130 = 4W + 10
Subtract 10 from each side
120 = 4W
Divide each side by 4
30 = W
If the width is 30, then the length is 35
30 + 30 + 35 + 35 = 130